# Optimizing Treatment Allocation in the Presence of Interference

### v1

@@ 5.1.2 Causal estimator
As already mentioned in Section 4, we will use a modified version of NetEst (Jiang & Sun, 2022) to estimate ITTEs. We opt for NetEst because of its strong empirical performance, solid theoretical foundation, and the availability of its source code. The architecture is shown in Figure 3. Inputs are given in blue  , (intermediate) outputs in yellow  , and the neural network layers in green  . NetEst uses an adversarial training approach to balance the representations $\mathbf{\phi}_{i}$ (Berrevoets et al., 2020). This is achieved by using two discriminators $d_{T}$ and $d_{Z}$ that predict the treatment $t_{i}$ and the peer exposure $z_{i}$, respectively. To balance the representations, the distributions $p(T_{i}|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}})$ and $p(Z_{i}|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},t_{i})$ are uniformed. More specifically, there should be no information in the hidden representation $\phi_{i}$ that allows to predict $t_{i}$ or $z_{i}$. To calculate $\phi_{i}$, a Graph Convolutional Network (GCN) (Kipf & Welling, 2016) and an MLP $e_{\phi}$ are used. $\phi_{i}$ is fed into the predictor $p_{Y}$ to make a prediction for the potential outcome ${y}_{i}$. The discriminators $d_{Z}$ and $d_{T}$, and the predictor $p_{Y}$ are MLPs with ReLU non-linearities that consist of 3 layers with each hidden layer consisting of 16 hidden nodes. $e_{\phi}$ has one layer with a ReLU non-linearity. The GCN consists of one convolutional layer (Jiang & Sun, 2022).
@@
These regularization losses are employed to uniform the distributions. Since $T_{i}$ is binary, we can use $0.5$ as the uniform distribution. For $Z_{i}$, $c_{i}$ is sampled uniformly at random from $\left[0,1\right]$ for each $i$ in each iteration. The outcome loss $\mathcal{L}_{Y}$ compares the prediction $\hat{y}_{i}$ with the actual outcomes $y_{i}$ using Binary Cross Entropy (BCE). $\mathcal{L}_{Y}$ is used to update the estimator $e_{Y}$. Finally, the combined loss is calculated:
@@
This loss is used to update the GCN and the encoder $e_{\phi}$ (Jiang & Sun, 2022).
The main modifications made to the original NetEst architecture are the change to a binary outcome $Y_{i}$ and the addition of the encoder $e_{\phi}$. The latter improves performance by separating $\mathbf{x}_{i}$ from $\mathbf{x}_{\mathcal{N}_{i}}$. Following Jiang & Sun (2022) we set $\alpha$ and $\gamma$ to 0.5. The Adam optimizer (Kingma & Ba, 2015) is employed to train the network. Furthermore, the neural network is trained using full-batch training. Code provided by Jiang & Sun (2022) was used as a starting point for both the DGP and NetEst implementation.
@@ 5.1.3 Optimization algorithms
To maximize TTE we use two heuristics: the greedy algorithm (OTAPI-GR) (Kempe et al., 2003) and the genetic algorithm (OTAPI-GA) (Bucur & Iacca, 2016). The greedy algorithm is an iterative approach where the allocation set is sequentially expanded until the budget $k$ is reached. In each iteration, the marginal gain in TTE associated with the addition of each remaining node to the allocation set is calculated. This marginal gain can be estimated as $\sum_{i\in\mathcal{N}}\hat{\omega}_{i}(t_{i}^{\prime},z_{i}^{\prime})-\sum_{i\in\mathcal{N}}\hat{\omega}_{i}(t_{i},z_{i})$ where $\mathbf{t}$ is the treatment vector from the previous iteration, and $\mathbf{t^{\prime}}$ is $\mathbf{t}$ with an additional node $j$ receiving treatment. The treatment vector $\mathbf{t^{\prime}}$ with the highest estimated TTE is chosen for that iteration.
The genetic algorithm (GA) iteratively improves a population of solutions, represented by treatment allocation vectors $\mathbf{t}$, by applying genetic operators such as crossover and mutation. First, a fitness function has to be defined. We employ predicted TTE. To incorporate the budget constraint, the fitness function returns zero when this constraint is violated. An initial set of solutions has to be generated. To converge more quickly, and to get better solutions, the degree (DEG) and single discount (SD) heuristic solutions are added to the initial population. The other solutions in the initial population are randomly generated. Thereafter, the best solutions are selected to be parents to create new solutions (children). These solutions differ from their parents by using crossover (a child is made up of the parts of its parents) and mutation (randomly changing small parts of the child). Thanks to this randomness, escaping from local optima is made possible. Finally, – to not lose any really good solutions – the best solutions from the previous iteration are kept unchanged and the next iteration starts (Gad, 2023; Bucur & Iacca, 2016). The genetic algorithm was implemented using the PyGAD library (Gad, 2023). The hyperparameters were tuned using the predicted TTE on the synthetic data with $\beta_{spillover}=0.3$. The hyperparameters are shown in Table 2. The number of generations is dependent on the budget $k$ because the algorithm converges faster for lower $k$.
@@ 5.1.4 Methods for comparison
  • TARNet: UM approach. TARNet (Shalit et al., 2017), a commonly used causal estimator, is used for ITE estimation. TARNet is trained on individual data only, not leveraging any network information. The $k$ entities with the highest ITE estimates are allocated treatment. We use the implementation for TARNet (Shalit et al., 2017) from Jiang & Sun (2022) and largely follow their setup. The model is trained using full-batch training with BCE loss and employs the Adam optimizer Kingma & Ba (2015). The number of hidden nodes in each layer is 16.
  • Upper Bound (UB): To get an idea of the upper bound on the evaluation metrics, we implement a method that leverages the true DGP, which simulates the performance of a perfect model $\mathcal{M}$. Since the allocation problem is – even with a perfect model – still a combinatorial problem, a heuristic has to be used to solve it. To this end, we use the greedy heuristic. Note that this means that the true upper bound may be better than obtained with this heuristic.
@@ 5.1.5 Evaluation metrics
We define Liftup in TTE for method $m$ and budget $k$ as
$Liftup_{m}(k)=\frac{TTE_{m}(k)}{TTE_{random}(k)}.$ (9)
This metric captures the relative increase in TTE of a method compared to randomly assigning an equal amount of treatments. Since there are many possible random assignments, we randomly sample 100 possible assignments for a budget $k$ and calculate the average TTE to estimate $TTE_{random}(k)$. 100 was chosen because it gave stable estimates for all $k$. Moreover, to get a more complete understanding of the results, the relative increase in sum of expected outcomes (RISEO) is defined as:
$RISEO_{m}(k)=\frac{\sum_{i\in\mathcal{N}}\mathbb{E}\bigl{[}Y_{i}(t_{i,m},z_{i,m})\bigr{]}}{\sum_{i\in\mathcal{N}}\mathbb{E}\bigl{[}Y_{i}(t_{i,random},z_{i,random})\bigr{]}}.$ (10)
@@ 5.1.6 Hyperparameter selection
We use grid search to find the best hyperparameters for TARNet and NetEst for each setup. To select the best model we use the validation Binary Cross Entropy (BCE) to compare the factual outcomes $y_{i}$ to the predicted outcomes $\hat{y}_{i}$. However, the model that is best at predicting the potential outcomes, is not necessarily the best at predicting causal effects. Although model selection in causal inference is still an unresolved issue, better approaches may exist (Curth & Van Der Schaar, 2023). However, given that these approaches have not yet been tested in a network context, we leave this for future research. The hyperparameter ranges for NetEst and TARNet are shown in Tables 3 and 4. The source code for all our experiments can be found in this repository: https://github.com/daan-caljon/OTAPI.
@@
Table 3: Hyperparameters NetEst
@@
Table 4: Hyperparameters TARNet
@@ 5.2.1 Influence of budget $k$
In LABEL:fig:semi-synthetic, we show the results for the test sets of the two semi-synthetic datasets BC and Flickr. We plot Liftup and RISEO against the budget $k$ and observe that OTAPI-GR and OTAPI-GA outperform (or have at least the same performance as) all other methods for all $k$. Furthermore, their performance is close to the Upper Bound, especially for the Flickr dataset. Note that OTAPI-GR and OTAPI-GA have an almost identical performance (making it hard to see the curves of OTAPI-GR in the figures). Moreover, "network-aware" methods (i.e., all methods except for TARNet) outperform the UM approach for relatively small budgets $k$. However, once a certain budget is surpassed, the TARNet method outperforms CELF, DEG, and SD for both datasets. This happens already for a small budget $k$ for the BC dataset. This pattern is expected since the degree distribution follows a power law (Barabási & Albert, 1999). This means that a relatively low number of nodes have a high degree. Intuitively, we would expect that treating these nodes should have a large impact on the network as a whole, which is confirmed by these results. This is further explored in Section 5.2.3. The UM approach (TARNet) fails to identify these high-impact entities. As $k$ increases, identifying nodes with a large marginal individual effect becomes more important than the spillover effects. OTAPI can capture both the spillover and marginal individual effects, which explains why it performs well for all $k$. Note that the peaks of the Liftup and RISEO curves have no direct implications in terms of decision-making. The optimal $k$ depends on the relative costs and benefits of the treatments and their effects, respectively (Verbeke et al., 2020).
@@ 5.2.3 Impact of the degree distribution
All networks investigated in this paper possess the scale-free property (Barabási & Albert, 1999). This means that the distribution of the degree follows a power law. As shown in LABEL:fig:degree_effect, a small number of nodes has a very high degree. Consequently, they will most likely have a large network effect. This may explain why the degree and single discount heuristics perform quite well for small $k$. To further investigate this hypothesis, we simulate a small dataset of 2,000 nodes using the Watts-Strogatz (Watts & Strogatz, 1998) random network model. Networks that follow this model do not possess the scale-free property. The degree distribution for the simulated data using the Watt-Strogatz model is shown in Figure LABEL:fig:Degree_WS. To make a fair comparison to the Barabási–Albert random networks, the parameters of the Watts-Strogatz random network were chosen to have the same average degree (4).
@@ 5.2.4 Allocation similarity between methods
To gain more insight into the similarities and differences between methods, we calculate a similarity matrix between the solutions for the Flickr test set at $k=5\%$, $k=20\%$, and $k=40\%$. Similarity is defined as the number of selected nodes that two methods have in common divided by the budget $k$. The similarity matrices are shown in LABEL:tab:sim_matrix. As expected, solutions with a larger Liftup (see LABEL:fig:Flickr_Liftup) are more similar to the UB solution. Furthermore, we observe that OTAPI-GR and OTAPI-GA yield an almost identical solution. The high similarity explains why the Liftup for both methods is almost identical. This similarity between the two methods persists for higher $k$. Furthermore, the ability to leverage network structure plays a significant role in finding solutions with high similarity to UB. This pattern changes as $k$ increases, then the similarity between UB and TAR, and OTAPI and TAR increases, while the similarity between UB and DEG, and UB and SD decreases. This is because, at some point when increasing $k$, the marginal individual effects gradually become more important than the spillover effects. If this analysis were to be extended to higher $k$, it would become increasingly difficult to extract insightful conclusions from the similarity matrix as a random solution would already have an expected similarity of $k\%$ with each solution. Consequently, $k=40\%$ is the highest budget we investigate using the similarity matrix.
@@ 6 Conclusion
We present OTAPI, a novel method to optimize treatment allocation in the presence of network interference. Currently, two alternative approaches, IM and UM, exist. OTAPI fills the gap between these methods by combining their strengths. Our method consists of two steps. First, a relational causal estimator is trained to accurately estimate treatment effects in the presence of interference. Second, because the problem of finding the optimal treatment allocation in a network setting is combinatorial in nature, optimization algorithms are leveraged. These algorithms intelligently employ the relational causal estimator for treatment allocation.
We have empirically demonstrated that OTAPI outperforms alternative treatment allocation methods. Moreover, the relative performance of our method is robust to both the amount of available treatments $k$ and the magnitude of the spillover effects. By contrast, the relative performance of IM and UM approaches is highly dependent on these parameters.
A limitation of our work is that the assumed causal structure might not fully capture reality, as it does not account for potential contagion effects, which means that the outcome of an entity at time $t$ can influence the outcome of another entity at time $t+1$ (Ogburn & VanderWeele, 2014). Moreover, we have followed existing literature (Forastiere et al., 2021; Jiang & Sun, 2022) by assuming that the exposure mapping $Z$ is defined as the ratio of neighbors treated. This assumption will likely be violated in practice. Consequently, the causal estimator that was used in this paper may become quite inaccurate. However, since OTAPI is agnostic to the relational causal estimator that is used, other estimators that do not need this assumption (e.g., Huang et al., 2023; Adhikari & Zheleva, 2023) can be seamlessly incorporated. Nevertheless, OTAPI’s performance using more complex relational causal estimators should be validated in future work. Furthermore, the estimation errors from the relational causal estimator might have an impact on the treatment allocation optimization in the second step of OTAPI. Therefore, an interesting avenue for future research is to investigate whether an end-to-end approach, where the treatment allocation is directly learned from the input data, is feasible for this problem. Finally, also incorporating potential costs and benefits of treatments and outcomes provides an interesting direction for future research to determine the optimal number of treatments $k^{*}$ in a data-driven manner (Verbeke et al., 2020, 2023).

### v2

@@ 5.1.2 Causal estimator
As already mentioned in Section 4, we will use a modified version of NetEst (Jiang and Sun, 2022) to estimate ITTEs. We opt for NetEst because of its strong empirical performance, solid theoretical foundation, and the availability of its source code. Note, however, that other relational causal estimators could also be used. The architecture is shown in Figure 3. Inputs are given in blue  , (intermediate) outputs in yellow  , and the neural network layers in green  . NetEst uses an adversarial training approach to balance the representations $\mathbf{\phi}_{i}$ (Berrevoets et al., 2020). This is achieved by using two discriminators $d_{T}$ and $d_{Z}$ that predict the treatment $t_{i}$ and the peer exposure $z_{i}$, respectively. To balance the representations, the distributions $p(T_{i}|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}})$ and $p(Z_{i}|\mathbf{x}_{i},\mathbf{x}_{\mathcal{N}_{i}},t_{i})$ are uniformed. More specifically, there should be no information in the hidden representation $\phi_{i}$ that allows to predict $t_{i}$ or $z_{i}$. To calculate $\phi_{i}$, a Graph Convolutional Network (GCN) (Kipf and Welling, 2016) and an MLP $e_{\phi}$ are used. $\phi_{i}$ is fed into the predictor $p_{Y}$ to make a prediction for the potential outcome ${y}_{i}$. The discriminators $d_{Z}$ and $d_{T}$, and the predictor $p_{Y}$ are MLPs with ReLU non-linearities that consist of 3 layers. $e_{\phi}$ has one layer with a ReLU non-linearity. The GCN consists of one convolutional layer (Jiang and Sun, 2022).
@@
These regularization losses are employed to uniform the distributions. Since $T_{i}$ is binary, we can use $0.5$ as the uniform distribution. For $Z_{i}$, $c_{i}$ is sampled uniformly at random from $\left[0,1\right]$ for each $i$ in each iteration. The outcome loss $\mathcal{L}_{Y}$ compares the prediction $\hat{y}_{i}$ with the actual outcomes $y_{i}$ using BCE. $\mathcal{L}_{Y}$ is used to update the estimator $e_{Y}$. Finally, the combined loss is calculated:
@@
This loss is used to update the GCN and the encoder $e_{\phi}$ (Jiang and Sun, 2022).
The main modifications made to the original NetEst architecture are the change to a binary outcome $Y_{i}$ and the addition of the encoder $e_{\phi}$. The latter improves performance by separating $\mathbf{x}_{i}$ from $\mathbf{x}_{\mathcal{N}_{i}}$. Following Jiang and Sun (2022) we set $\alpha$ and $\gamma$ to 0.5. The Adam optimizer (Kingma and Ba, 2015) is employed to train the network. Furthermore, the neural network is trained using full-batch training. Code provided by Jiang and Sun (2022) was used as a starting point for both the DGP and NetEst implementation.
@@ 5.1.3 Optimization algorithms
To maximize TTE we use two heuristics: the greedy algorithm (OTAPI-GR) and the genetic algorithm (OTAPI-GA).
The greedy algorithm is an iterative approach where the allocation set is sequentially expanded until the budget $k$ is reached. In each iteration, the marginal gain in TTE associated with the addition of each remaining node to the allocation set is calculated. This marginal gain can be estimated as $\sum_{i\in\mathcal{N}}\hat{\omega}_{i}(t_{i}^{\prime},z_{i}^{\prime})-\sum_{i\in\mathcal{N}}\hat{\omega}_{i}(t_{i},z_{i})$ where $\mathbf{t}$ is the treatment vector from the previous iteration, and $\mathbf{t^{\prime}}$ is $\mathbf{t}$ with an additional node $j$ receiving treatment. The treatment vector $\mathbf{t^{\prime}}$ with the highest estimated TTE is chosen for that iteration.
The Genetic Algorithm (GA) iteratively improves a population of solutions, represented by treatment allocation vectors $\mathbf{t}$, by applying genetic operators such as crossover and mutation. The fitness function used by the GA is the TTE obtained by using the trained causal model. To incorporate the budget constraint, the fitness function returns zero when this constraint is violated. An initial set of solutions has to be generated. To converge more quickly, and to get better solutions, the degree and single discount heuristic solutions are added to the initial population. The other solutions in the initial population are randomly generated. Thereafter, the best solutions are selected to be parents to create new solutions (children). These solutions differ from their parents by using crossover (a child is made up of the parts of its parents) and mutation (randomly changing small parts of the child). Thanks to this randomness, escaping from local optima is made possible. Finally, to avoid losing any particularly good solutions, the best solutions from the previous iteration are kept unchanged, and the next iteration begins (Gad, 2023; Bucur and Iacca, 2016). The GA was implemented using the PyGAD library (Gad, 2023). The hyperparameters were tuned using the predicted TTE on the synthetic data with $\beta_{spillover}=0.3$. The hyperparameters are shown in Table 2. The number of generations is dependent on the budget $k$ because the algorithm converges faster for lower $k$.
@@ 5.1.4 Methods for comparison
  • TARNet: UM approach. TARNet (Shalit et al., 2017), a commonly used causal estimator, is used for ITE estimation. TARNet is trained on individual data only, not leveraging any network information. The $k$ entities with the highest ITE estimates are allocated treatment. We use the implementation for TARNet from Jiang and Sun (2022) and largely follow their setup. The model is trained using full-batch training with BCE loss and employs the Adam optimizer Kingma and Ba (2015).
  • Oracle Greedy (OG): To get an idea of the best possible performance on the evaluation metrics, we implement a method that leverages the true DGP, which simulates the performance of a perfect model $\mathcal{M}$. Since the allocation problem is—even with a perfect model—still an NP-hard problem, a heuristic is used to solve it. To this end, we use the greedy heuristic. While there is no guarantee that this solution is close to the true optimal treatment allocation, it serves as a useful baseline to provide insight into the performance of the different methods.
@@ 5.1.5 Evaluation metrics
We define Liftup in TTE for method $m$ and budget $k$ as:
$\text{Liftup}_{m}(k)=\frac{\text{TTE}_{m}(k)}{\text{TTE}_{random}(k)}.$ (7)
This metric captures the relative increase in TTE of a method compared to randomly assigning an equal amount of treatments. Since there are many possible random assignments, we randomly sample 100 possible assignments for a budget $k$ and calculate the average TTE to estimate $\text{TTE}_{random}(k)$. 100 was chosen because it gave stable estimates for all $k$. Moreover, to get a more complete understanding of the results, the relative increase in sum of expected outcomes (RISEO) is defined as:
$\text{RISEO}_{m}(k)=\frac{\sum_{i\in\mathcal{N}}\mathbb{E}\bigl{[}Y_{i}(t_{i,m},z_{i,m})\bigr{]}}{\sum_{i\in\mathcal{N}}\mathbb{E}\bigl{[}Y_{i}(t_{i,random},z_{i,random})\bigr{]}}.$ (8)
@@ 5.1.6 Hyperparameter selection
We use grid search to find the best hyperparameters for TARNet and NetEst for each setup. To select the best model, we use the validation BCE to compare the factual outcomes $y_{i}$ to the predicted outcomes $\hat{y}_{i}$. However, the model that is best at predicting the potential outcomes, is not necessarily the best at predicting causal effects. Although model selection in causal inference is still an unresolved issue, better approaches may exist (Curth and Van Der Schaar, 2023). However, given that these approaches have not yet been tested in a network context, we leave this for future research. The hyperparameter ranges for NetEst and TARNet are shown in Tables 3 and 4. Given the high feature dimensionality in the Enron dataset, we also select the best hidden layer dimension using a grid search over $\{8,16,32\}$ for both NetEst and TARNet. For all other datasets, the hidden layer dimension is set to 16. The source code for all our experiments can be found in this repository: https://github.com/daan-caljon/OTAPI.
@@
Table 3: Hyperparameters NetEst.
@@
Table 4: Hyperparameters TARNet.
@@ 5.2.1 Influence of budget $k$
In LABEL:fig:semi-synthetic, we show the results for the test sets of the two semi-synthetic datasets BC and Flickr. We plot Liftup and RISEO against the budget $k$ and observe that OTAPI-GR and OTAPI-GA outperform (or have at least the same performance as) all other methods for all $k$. Furthermore, their performance is close to the OG solution, especially for the Flickr dataset. Note that OTAPI-GR and OTAPI-GA have an almost identical performance (making it hard to see the curves of OTAPI-GR in the figures). Moreover, “network-aware" methods (i.e., all methods except for TARNet) outperform the UM approach for relatively small budgets $k$. However, once a certain budget is surpassed, the TARNet method outperforms CELF, DEG, and SD for both datasets. This happens already for a small budget $k$ for the BC dataset. This pattern is expected since the degree distribution follows a power law (Barabási and Albert, 1999). This means that a relatively low number of nodes have a high degree. Intuitively, we would expect that treating these nodes should have a large impact on the network as a whole, which is confirmed by these results. This is further explored in Section 5.2.3. The UM approach (TARNet) fails to identify these high-impact entities. As $k$ increases, identifying nodes with a large MITE becomes more important than the spillover effects. OTAPI can capture both the spillover effects and MITEs, which explains why it performs well for all $k$. Note that the peaks of the Liftup and RISEO curves have no direct implications in terms of decision-making. The optimal $k$ depends on the relative costs and benefits of the treatments and their effects, respectively (Verbeke et al., 2020).
@@ 5.2.3 Impact of the degree distribution
All networks investigated in this paper possess the scale-free property (Barabási and Albert, 1999). This means that the distribution of the degree follows a power law. As shown in LABEL:fig:degree_effect, a small number of nodes has a very high degree. Consequently, they will most likely have a large network effect. This may explain why DEG and SD perform quite well for small $k$. To further investigate this hypothesis, we simulate a small dataset of 2,000 nodes using the Watts-Strogatz (Watts and Strogatz, 1998) random network model. Networks that follow this model do not possess the scale-free property. The degree distribution for the simulated data using the Watt-Strogatz model is shown in Figure LABEL:fig:Degree_WS. To make a fair comparison to the Barabási–Albert random networks, the parameters of the Watts-Strogatz random network were chosen to have the same average degree (4).
@@ 5.2.4 Allocation similarity between methods
To gain more insight into the similarities and differences between methods, we calculate a similarity matrix between the solutions for the Flickr test set at $k=5\%$, $k=20\%$, and $k=40\%$. Similarity is defined as the number of selected nodes that two methods have in common divided by the budget $k$. The similarity matrices are shown in LABEL:tab:sim_matrix. As expected, solutions with a larger Liftup (see LABEL:fig:Flickr_Liftup) are more similar to the OG solution. Furthermore, we observe that OTAPI-GR and OTAPI-GA yield an almost identical solution. The high similarity explains why the Liftup for both methods is almost identical. This similarity between the two methods persists for higher $k$. Furthermore, the ability to leverage network structure plays a significant role in finding solutions with high similarity to OG. This pattern changes as $k$ increases, then the similarity between OG and TARNet, and OTAPI and TARNet increases, while the similarity between OG and DEG, and OG and SD decreases. This is because, at some point when increasing $k$, the MITEs gradually become more important than the spillover effects. If this analysis were to be extended to higher $k$, it would become increasingly difficult to extract insightful conclusions from the similarity matrix as a random solution would already have an expected similarity of $k\%$ with each solution. Consequently, $k=40\%$ is the highest budget we investigate using the similarity matrix.
5.2.5 Potential influence of dataset size and feature dimensionality
Given the higher dimensionality and larger size of the Enron dataset, we investigate whether the relative performance of the different methods remains consistent. Due to the large number of nodes, OG and CELF are computationally too expensive and are therefore omitted (see Section 5.2.6). The results, shown in LABEL:fig:Enron_figs, are similar to those presented for the other datasets. OTAPI achieves the best performance, closely followed by DEG and SD for the lowest budgets. As the budget increases, the performance of DEG and SD deteriorates relative to OTAPI, and at $k=5\%$, their performance is comparable to that of TARNet. This experiment suggests that as both the feature dimensionality and the number of nodes increase, the relative performance of the different methods remains consistent.
5.2.6 Empirical analysis of runtime
In LABEL:fig:runtime_figs, the runtimes of the different methods are shown across a range of simulated dataset sizes for different budgets $k$. Note that both axes in the figures are on a logarithmic scale. All experiments were conducted on a machine with an Intel Xeon Gold 6140 CPU @ 2.3 GHz, 45 GiB of RAM, and an NVIDIA P100 GPU. If an algorithm exceeds the maximum runtime of 48 hours, it is terminated and its runtime is excluded from the figure. The figures show that the methods based on a greedy approach (OTAPI-GR, CELF, and OG) exhibit a substantial increase in runtime as the number of nodes grows. The runtime of the iterative SD approach also increases with dataset size. TARNet and DEG have the lowest runtimes, as they only require calculating a metric (degree or estimated TTE) and then sorting all instances according to this metric. The runtime of OTAPI-GA increases only slightly with dataset size.
@@ 6 Conclusion
We present OTAPI, a novel method to optimize treatment allocation in the presence of network interference. Currently, two alternative approaches, IM and UM, exist. OTAPI fills the gap between these methods by combining their strengths. Our method consists of two steps. First, a relational causal estimator is trained on observational data to accurately estimate treatment effects in the presence of interference. Second, the combinatorial problem of finding the optimal treatment allocation $\mathbf{t}^{*}$ for a potentially unseen network is solved by using an optimization algorithm that intelligently employs treatment effect estimates from the trained relational causal estimator from the first step. OTAPI is agnostic to the specific method used in each step; that is, any relational causal estimator and any optimization algorithm from the classic IM literature can be used.
We have empirically demonstrated that OTAPI outperforms alternative treatment allocation methods. Moreover, the relative performance of our method is robust to both the amount of available treatments $k$ and the magnitude of the spillover effects. By contrast, the relative performance of IM and UM approaches is highly dependent on these parameters. Additional experiments showed that OTAPI maintained its superior relative performance across varying degree distributions, dataset sizes, and feature dimensionalities, outperforming the baselines in every setting considered.
A limitation of our work is that the assumed causal structure might not fully capture reality, as it does not account for potential contagion effects, which means that the outcome of an entity at time $t$ can influence the outcome of another entity at time $t+1$ (Ogburn and VanderWeele, 2014). Moreover, we have followed existing literature (Forastiere et al., 2021; Jiang and Sun, 2022) by assuming that the exposure mapping $Z$ is defined as the ratio of neighbors treated. This assumption will likely be violated in practice. Consequently, the causal estimator that was used in this paper may become quite inaccurate. However, since OTAPI is agnostic to the relational causal estimator that is used, other estimators that do not need this assumption (e.g., Huang et al., 2023; Adhikari and Zheleva, 2023) can be seamlessly incorporated. Nevertheless, OTAPI’s performance using more complex relational causal estimators should be validated in future work. Furthermore, the estimation errors from the relational causal estimator might have an impact on the treatment allocation optimization in the second step of OTAPI. Therefore, an interesting avenue for future research is to investigate whether an end-to-end approach, where the treatment allocation is directly learned from the input data, is feasible for this problem. Finally, also incorporating potential costs and benefits of treatments and outcomes provides an interesting direction for future research to determine the optimal number of treatments $k^{*}$ in a data-driven manner (Verbeke et al., 2020, 2023).